Solve the system of equations by substitution: y= xsquared+2x-2 and y=3x+4
thanx

Answer by meghanjax13(2)

y=x^2+2x-2
y=3x+4
substitute 3x+4 for y in the first equation.
3x+4=x^2+2x-2
Then solve.
0=x^2-x-6
Factor.
x=3
Substitute x back into the equation to find y.
y=3(3)+4
y=13
Check by substituting x and y into the first equation to make sure it works.
y=x^2+2x-2
13=3^2+2(3)-2
13=13
Hope I could help.